# On the non-Markovian quantum control dynamics

### v3

@@ II-A Single multi-level atom coupled with the cavity
where $\chi_{n}$ is a constant as in Eq. (8) and Appendix A. Eq. (9a) represents that the atomic state excited at the $n$-th energy level can be acquired from a lower energy level by absorbing one photon from the cavity or from a higher energy level by emitting one photon into the cavity, which are represented by the first four components on the RHS, and the following components represent the non-Markovian interaction between the atom and environment, which can make the atom decay to a lower energy level. Eq. (9b) and Eq. (9c) represent the interface between the atom and the cavity via the emitting and absorbing processes of a photon. This process is influenced by the detuning between the atom and the cavity (i.e., the first component of the RHS of Eq. (9b) and Eq. (9c)), the coupling strengths $\tilde{g}_{n}$, and the non-Markovian decaying to the environment represented by the last component of the RHS of Eq. (9b) and Eq. (9c), respectively. Eq. (9d) represents how the number of photons in the cavity is influenced by the coupling between the atom and cavity. The nonlinear Eq. (9e) represents the non-Markovian decaying rate of the atom to the environment, which is derived in detail in Appendix A. Additionally, the atom populations are normalized as $\langle\sigma_{00}\rangle+\sum_{n=1}^{N-1}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle=1$, where $\langle\sigma_{00}\rangle$ represents the population that the atom is at the ground state and its dynamics is governed as
$\frac{\mathrm{d}}{\mathrm{d}t}\langle\sigma_{00}\rangle=$ $\left[F_{1}(t)+F_{1}^{*}(t)\right]\kappa_{1}\langle\sigma^{+}_{1}\sigma^{-}_{1}\rangle$ (10)
$+i\tilde{g}_{1}\langle\sigma^{+}_{1}a\rangle-i\tilde{g}_{1}^{*}\langle\sigma^{-}_{1}a^{{\dagger}}\rangle.$
We define the state vector
@@
$A_{n}=$ $-\kappa_{n}{\rm{diag}}\left[(F_{n}+F_{n}^{*}),F_{n}^{*},F_{n}\right]$ (13)
$+\begin{bmatrix}0&-i\tilde{g}_{n}&i\tilde{g}_{n}^{*}\\ -i\tilde{g}_{n}^{*}&-i\Delta_{n}&0\\ i\tilde{g}_{n}&0&i\Delta_{n}\end{bmatrix},$
@@
For the basic control model of this paper in Eq. (12), which is a time-varying linear equation, the state vector $X(t)$ in Eq. (11) is made up of the atomic state and photonic state in the cavity. The target of this paper is to study the condition for the occurrence of non-Markovian interactions between quantum system and environment evaluated by the nonlinear dynamics in Eq. (9e), and how the stability of time-varying atomic and photonic dynamics stability in Eq. (12) can be influenced by the non-Markovian process and additional control methods in the following.
@@ II-B Non-Markovian parameter dynamics
II-B1 $\Omega=\tilde{\omega}_{n}$ and $04\kappa_{n}S_{n}-Q_{n}^{2}<$ 0
@@
II-B2 $\Omega=\tilde{\omega}_{n}$ and $04\kappa_{n}S_{n}-Q_{n}^{2}>$ 0
@@
$F_{n}=$ $-\frac{Q_{n}}{2\kappa_{n}}+\sqrt{\frac{S_{n}}{\kappa_{n}}-\left(\frac{Q_{n}}{2\kappa_{n}}\right)^{2}}$ (20)
×$\times\tan\left[\sqrt{\frac{S_{n}}{\kappa_{n}}-\left(\frac{Q_{n}}{2\kappa_{n}}\right)^{2}}\left(\kappa_{n}t+C\right)\right],$
@@
II-B3 $\Omega=\tilde{\omega}_{n}$ and $04\kappa_{n}S_{n}-Q_{n}^{2}=$ 0
@@
When $04\kappa_{n}S_{n}-Q_{n}^{2}>$ 0, $\lim_{t\rightarrow\infty}F_{n}(t)$ does not exist because $\tan\left[\sqrt{\frac{S_{n}}{\kappa_{n}}-\left(\frac{Q_{n}}{2\kappa_{n}}\right)^{2}}\left(\kappa_{n}t+C\right)\right]$ is infinite when
$\sqrt{\frac{S_{n}}{\kappa_{n}}-\left(\frac{Q_{n}}{2\kappa_{n}}\right)^{2}}\left(\kappa_{n}t+C\right)=l\pi+\pi/2,$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}X_{F}(t)=\begin{bmatrix}\kappa_{n}R_{n}-\gamma&-\kappa_{n}I_{n}\\ \kappa_{n}I_{n}&\kappa_{n}R_{n}-\gamma\end{bmatrix}\begin{bmatrix}R_{n}\\ I_{n}\end{bmatrix}$ (24)
$+\begin{bmatrix}0&\left(\Omega-\tilde{\omega}_{n}\right)\\ -\left(\Omega-\tilde{\omega}_{n}\right)&0\end{bmatrix}\begin{bmatrix}R_{n}(t)\\ I_{n}(t)\end{bmatrix}+\begin{bmatrix}\frac{\gamma\chi_{n}}{2}\\ 0\end{bmatrix}$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\delta X_{F}(t)^{\rm T}\delta X_{F}(t)\right)$ (26)
$=$ $2\delta X_{F}(t)^{\rm T}\frac{\partial\textbf{f}}{\partial X_{F}}\delta X_{F}(t)+2\textbf{f}_{\Omega}\delta X_{F}(t)^{\rm T}\delta X_{F}(t).$
@@
For the nonlinear system in Eq. (23), we regard $u=\Omega-\tilde{\omega}_{n}$ as an unknown input, which can be regarded as a constant perturbation or random unknown uncertainty due to the fact that the environment cannot be precisely modeled. Then
$\frac{\mathrm{d}}{\mathrm{d}t}X_{F}(t)=\textbf{f}\left(X_{F}(t)\right)+\begin{bmatrix}0&u\\ -u&0\end{bmatrix}X_{F}(t),$ (28)
and we define the output of the system as
@@
We rewrite Eq. (24) as
$\frac{\mathrm{d}}{\mathrm{d}t}X_{F}(t)$ $=\left(\mathbf{A}\left(X_{F}(t)\right)+\mathbf{B}u\right)X_{F}(t)+\begin{bmatrix}\gamma\chi_{n}/2\\ 0\end{bmatrix},$ (30)
where $\textbf{f}\left(X_{F}(t)\right)=\mathbf{A}\left(X_{F}(t)\right)X_{F}(t)+\begin{bmatrix}\gamma\chi_{n}/2\\ 0\end{bmatrix}$, $\mathbf{B}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$, and $\mathbf{A}$ is the first matrix on the RHS of Eq. (24).
@@
$\dot{V}(X_{F})=$ $\textbf{f}^{\rm T}\left(X_{F}(t)\right)\left[\frac{\partial\textbf{f}^{\rm T}}{\partial X_{F}}\mathbf{M}\left(X_{F}\right)\right.$ (31)
$\left.+\mathbf{M}\left(X_{F}\right)\frac{\partial\textbf{f}}{\partial X_{F}}+\dot{\mathbf{M}}\right]\textbf{f}\left(X_{F}(t)\right).$
@@ II-B4 Example
Take the parameter settings in Section II-B3) as an example. When $\Omega=\tilde{\omega}_{n}$ and $4\kappa_{n}S_{n}=Q_{n}^{2}$, $I_{n}(t)\equiv 0$, $Q_{n}=-\gamma$, $S_{n}=\frac{\gamma^{2}}{4\kappa_{n}}$, $R_{n}(t)=\frac{\gamma}{2\kappa_{n}}-\frac{\gamma}{\gamma\kappa_{n}t+2\kappa_{n}}$, then $\kappa_{n}R_{n}(t)-\gamma<0$ for arbitrary $t$ and $\textbf{J}\left(X_{F}(t)\right)$ is negative definite. If we take $\mathbf{M}=\mathbf{I}$ as the identity matrix for simplification in Eq. (31), then $\dot{V}(X_{F})<0$ and Eq. (23) is globally stable in this parameter setting.
@@
When $\Omega=\tilde{\omega}_{n}$ in Eq. (24) and $\mathbf{M}=\mathbf{I}$ in Eq. (31), there exists $\alpha_{V}>0$ such that the parameters in Eq. (24) satisfy $02\kappa_{n}\alpha_{V}+\gamma\chi_{n}-2\gamma R_{n}<$ 0. In this case, there exists an invariant set of $X_{F}$ represented as $\Omega_{\alpha}=\left\{X_{F}\in\textbf{R}^{2}:V\leq\alpha_{V}\right\}$ contained in the region satisfying $\dot{V}=0$, and $\alpha_{V}\leq\gamma(\gamma-\chi_{n}\kappa_{n})/\kappa_{n}^{2}$.
@@
Assume initially $X_{F}\in\Omega_{\alpha}$ and $\Gamma_{F}$ represents an invariant set of $X_{F}$ satisfying that $\dot{V}(X_{F})=0$. When $02\kappa_{n}\alpha_{V}+\gamma\chi_{n}-2\gamma R_{n}<$ 0, $\dot{V}<0$ is satisfied according to Eqs. (31), (23a) with $\Omega=\tilde{\omega}_{n}$, and $V(x)\geq 0$ by its definition. Then there exists a positive value $v_{l}$ that $\lim_{t\rightarrow\infty}V(X_{F})=v_{l}\geq 0$ and $v_{l}<\alpha_{V}$. Thus $V(X_{F})$ is always within the invariant set. Denote the solutions of $R_{n}$ satisfying $\dot{V}=0$ as $r_{1}$ and $r_{2}$, then $\alpha_{V}\leq r_{1}^{2}+r_{2}^{2}=\gamma(\gamma-\chi_{n}\kappa_{n})/\kappa_{n}^{2}$. ∎
@@
Taking the circumstance $N=2$ as an example, the real and imaginary parts of $F_{1}(t)$ are compared in Fig. 2, where we take $\chi_{1}=1$, $\gamma=2$GHz [53], $\kappa_{1}=1$GHz, $\Omega=50$GHz and $\mathrm{d}t=0.01$ns. The comparisons in Figs. 2(a) and (b) show that, the diverse between $\Omega$ and $\tilde{\omega}_{1}$ can induce oscillations of $R_{1}(t)$ and $I_{1}(t)$, and this can influence the non-Markovian dynamics by via the time-varying decaying rates in Eq. (9).
@@
II-C Nonlinear dynamics with uncertain input
@@
According to Eqs. (18)-(21), when $\Omega=\tilde{\omega}_{n}$, $I_{n}(t)\equiv 0$, and we denote $\lim_{t\rightarrow\infty}R_{n}(t)=\bar{R}_{n}$. Define a new state vector $\tilde{X}_{F}(t)=\left[R_{n}(t)-\bar{R}_{n},I_{n}(t)\right]^{\rm T}\triangleq\left[\tilde{R}_{n}(t),I_{n}(t)\right]^{\rm T}$ and $\lim_{t\rightarrow\infty}\tilde{R}_{n}(t)=\left[0,0\right]^{\rm T}\triangleq\tilde{X}_{F}^{*}$ when Assumption 1 is satisfied and $\Omega=\tilde{\omega}_{n}$. Then
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{R}_{n}(t)=\kappa_{n}\left[\left(\tilde{R}_{n}(t)+\bar{I}_{n}\right)^{2}-I_{n}^{2}(t)\right]$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}-\gamma\left(\tilde{R}_{n}(t)+\bar{R}_{n}\right)+\left(\Omega-\tilde{\omega}_{n}\right)I_{n}(t)+\frac{\gamma\chi_{n}}{2},$ (33a)
$\frac{\mathrm{d}}{\mathrm{d}t}I_{n}(t)=2\kappa_{n}\left(\tilde{R}_{n}(t)+\bar{R}_{n}\right)I_{n}(t)$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}-\left(\Omega-\tilde{\omega}_{n}\right)\left(\tilde{R}_{n}(t)+\bar{R}_{n}\right)-\gamma I_{n}(t),$ (33b)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{X}_{F}(t)=$ $\tilde{\textbf{f}}\left(\tilde{X}_{F}(t)\right)+\begin{bmatrix}\frac{\gamma\chi_{n}}{2}+\kappa_{n}\bar{R}_{n}^{2}-\gamma\bar{R}_{n}\\ \left(\tilde{\omega}_{n}-\Omega\right)\bar{R}_{n}\end{bmatrix}$ (34)
$+\begin{bmatrix}0&\Omega-\tilde{\omega}_{n}\\ \tilde{\omega}_{n}-\Omega&0\end{bmatrix}\tilde{X}_{F}(t),$
@@
When $\tilde{\omega}_{n}=\Omega$, by Theorem 1, the quantum dynamics approaches Markovian behavior as $t\rightarrow\infty,$ and the nonlinear dynamics $\frac{\mathrm{d}}{\mathrm{d}t}\tilde{X}_{F}(t)=\tilde{\textbf{f}}\left(\tilde{X}_{F}(t)\right)$ is uniformly and asymptotically stable around its equilibrium $\tilde{X}_{F}=0$. When $\left|\tilde{\omega}_{n}-\Omega\right|\leq\epsilon$ and $\gamma\chi$ is finite, Proposition 6 in Appendix C ensures that $\tilde{R}_{n}(t)$ and $I_{n}(t)$ in Eq. (33) remain stable. Then the trajectory of $X(t)$ in Eq. (9) converges to be Markovian as $t\rightarrow\infty$. ∎
@@
Now we take $N=3$ as an example for numerical simulations of the above non-Markovian dynamics around the parameter settings in Refs. [54, 55]. As shown in Fig. 3, we take $\omega_{1}=37.7$GHz, $\omega_{2}=75.3$GHz, $\Omega=50$GHz, $\Delta_{1}=-20$MHz, $\Delta_{2}=-70$MHz, $g_{1}=g_{2}=20$MHz, $\chi_{1}=\chi_{2}=1$, $\gamma=10$GHz, $\kappa_{1}=\kappa_{2}=0.31$GHz, and the time step for simulation equals $0.01$ns. In the simulations for Markovian circumstance, we take $F_{n}$ with $n=1,2$ as constants that equal the steady values of those in the non-Markovian simulations.
Figure 3: Non-Markovian and Markovian dynamics of the quantum system with one three-level atom in a cavity.

### v4

@@ II-A Single multi-level atom coupled to the cavity
where $n=1,2,\dots,N-1$. Eq. (9a) represents that the atomic state excited at the $n$-th energy level can be acquired from a lower energy level by absorbing one photon from the cavity or from a higher energy level by emitting one photon into the cavity, which are represented by the first four components on the RHS, and the following components represent the non-Markovian interaction between the atom and environment, which can make the atom decay to a lower energy level. Eq. (9b) and Eq. (9c) represent the interface between the atom and the cavity via the emitting and absorbing processes of a photon. This process is influenced by the detuning between the atom and the cavity (i.e., the first component of the RHS of Eq. (9b) and Eq. (9c)), the coupling strengths $\tilde{g}_{n}$, and the non-Markovian decay to the environment represented by the last component of the RHS of Eq. (9b) and Eq. (9c), respectively. Eq. (9d) represents how the number of photons in the cavity is influenced by the coupling between the atom and cavity. The nonlinear Eq. (9e) represents the non-Markovian decay rate of the atom to the environment, which is derived in detail in Appendix A. Additionally, the atom populations are normalized as $\langle\sigma_{00}\rangle+\sum_{n=1}^{N-1}\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle=1$, where $\langle\sigma_{00}\rangle$ represents the population that the atom is at the ground state and $\langle\dot{\sigma}_{00}\rangle=\left[F_{1}+F_{1}^{*}\right]\kappa_{1}\langle\sigma^{+}_{1}\sigma^{-}_{1}\rangle+i\tilde{g}_{1}\langle\sigma^{+}_{1}a\rangle-i\tilde{g}_{1}^{*}\langle\sigma^{-}_{1}a^{{\dagger}}\rangle$.
We define the complex-valued state vector
@@
$A_{n}=$ $-\kappa_{n}{\rm{diag}}\left[(F_{n}+F_{n}^{*}),F_{n}^{*},F_{n}\right]+\begin{bmatrix}0&-i\tilde{g}_{n}&i\tilde{g}_{n}^{*}\\ -i\tilde{g}_{n}^{*}&-i\Delta_{n}&0\\ i\tilde{g}_{n}&0&i\Delta_{n}\end{bmatrix}$, (12)
@@
Remark 2.
By deriving the dynamics for the mean values of the operators representing atom’s populations, atom-cavity interactions, and the number of photons in the cavity, the non-Markovian equations (4,6) containing integral interactions with the environment can be analyzed from the perspective of linear time-varying dynamics in Eq. (11).
For the basic control model of this paper in Eq. (11), the parameter matrices are time-varying because the decay rates of the multi-level atom to the environment $F_{n}(t)$ are determined by time-varying integrations with memory effects. The target of this paper is to study the non-Markovian interactions between quantum system and environment evaluated by the nonlinear dynamics in Eq. (9e), and how the stability of time-varying atomic and photonic dynamics in Eq. (11) can be influenced by the non-Markovian process and additional control methods.
@@ II-B Non-Markovian parameter dynamics
II-B1 $\Omega=\tilde{\omega}_{n}$ and $4\kappa_{n}S_{n}-Q_{n}^{2}<0$
@@
II-B2 $\Omega=\tilde{\omega}_{n}$ and $4\kappa_{n}S_{n}-Q_{n}^{2}>0$
@@
$F_{n}=$ $-\frac{Q_{n}}{2\kappa_{n}}+\sqrt{\frac{S_{n}}{\kappa_{n}}-\left(\frac{Q_{n}}{2\kappa_{n}}\right)^{2}}\tan\left[\sqrt{\frac{S_{n}}{\kappa_{n}}-\left(\frac{Q_{n}}{2\kappa_{n}}\right)^{2}}\left(\kappa_{n}t+C\right)\right],$ (19)
@@
II-B3 $\Omega=\tilde{\omega}_{n}$ and $4\kappa_{n}S_{n}-Q_{n}^{2}=0$
@@
When $4\kappa_{n}S_{n}-Q_{n}^{2}>0$, $\lim_{t\rightarrow\infty}F_{n}(t)$ does not exist because $\tan\left[\sqrt{\frac{S_{n}}{\kappa_{n}}-\left(\frac{Q_{n}}{2\kappa_{n}}\right)^{2}}\left(\kappa_{n}t+C\right)\right]$ is infinite when
$\sqrt{\frac{S_{n}}{\kappa_{n}}-\left(\frac{Q_{n}}{2\kappa_{n}}\right)^{2}}\left(\kappa_{n}t+C\right)=l\pi+\frac{\pi}{2},$
@@
$\frac{\mathrm{d}}{\mathrm{d}t}X_{F}(t)$ $=\begin{bmatrix}\kappa_{n}R_{n}-\gamma&-\kappa_{n}I_{n}\\ \kappa_{n}I_{n}&\kappa_{n}R_{n}-\gamma\end{bmatrix}\begin{bmatrix}R_{n}\\ I_{n}\end{bmatrix}+\begin{bmatrix}0&\left(\Omega-\tilde{\omega}_{n}\right)\\ -\left(\Omega-\tilde{\omega}_{n}\right)&0\end{bmatrix}\begin{bmatrix}R_{n}(t)\\ I_{n}(t)\end{bmatrix}+\begin{bmatrix}\frac{\gamma\chi_{n}}{2}\\ 0\end{bmatrix}$ (23)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\delta X_{F}(t)^{\rm T}\delta X_{F}(t)\right)=2\delta X_{F}(t)^{\rm T}\frac{\partial\textbf{f}}{\partial X_{F}}\delta X_{F}(t)+2\textbf{f}_{\Omega}\delta X_{F}(t)^{\rm T}\delta X_{F}(t).$ (25)
@@
For the nonlinear system in Eq. (22), we regard $u=\Omega-\tilde{\omega}_{n}$ as an unknown input, representing a constant perturbation or random unknown uncertainty due to the fact that the environment cannot be precisely modeled. Then
$\frac{\mathrm{d}X_{F}(t)}{\mathrm{d}t}$ $=\textbf{f}\left(X_{F}(t)\right)+\begin{bmatrix}0&u\\ -u&0\end{bmatrix}X_{F}(t)$ (27)
$=\left(\mathbf{A}\left(X_{F}(t)\right)+\mathbf{B}u\right)X_{F}(t)+\begin{bmatrix}\gamma\chi_{n}/2\\ 0\end{bmatrix}$,
where $\textbf{f}\left(X_{F}(t)\right)=\mathbf{A}\left(X_{F}(t)\right)X_{F}(t)+\begin{bmatrix}\gamma\chi_{n}/2\\ 0\end{bmatrix}$, $\mathbf{A}$ is the first matrix on the RHS of Eq. (23), $\mathbf{B}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$, and we define the output of the system as
@@
$\dot{V}(X_{F})=$ $\textbf{f}^{\rm T}\left(X_{F}(t)\right)\left[\frac{\partial\textbf{f}^{\rm T}}{\partial X_{F}}\mathbf{M}\left(X_{F}\right)+\mathbf{M}\left(X_{F}\right)\frac{\partial\textbf{f}}{\partial X_{F}}+\dot{\mathbf{M}}\right]\textbf{f}\left(X_{F}(t)\right).$ (29)
@@ II-B4 Example
Take the parameter settings in Section II-B3) as an example. When $\Omega=\tilde{\omega}_{n}$ and $4\kappa_{n}S_{n}=Q_{n}^{2}$, $I_{n}(t)\equiv 0$, $Q_{n}=-\gamma$, $S_{n}=\gamma^{2}/4\kappa_{n}$, $R_{n}(t)=\gamma/2\kappa_{n}-\gamma/\left(\gamma\kappa_{n}t+2\kappa_{n}\right)$, then $\kappa_{n}R_{n}(t)-\gamma<0$ for arbitrary $t$ and $\textbf{J}\left(X_{F}(t)\right)$ is negative definite. If we take $\mathbf{M}=\mathbf{I}$ as the identity matrix for simplification in Eq. (29), then $\dot{V}(X_{F})<0$ and Eq. (22) is globally stable in this parameter setting.
@@
When $\Omega=\tilde{\omega}_{n}$ in Eq. (23) and $\mathbf{M}=\mathbf{I}$ in Eq. (29), there exists $\alpha_{V}>0$ such that the parameters in Eq. (23) satisfy $2\kappa_{n}\alpha_{V}+\gamma\chi_{n}-2\gamma R_{n}<0$. In this case, there exists an invariant set of $X_{F}$ represented as $\Omega_{\alpha}=\left\{X_{F}\in\textbf{R}^{2}:V\leq\alpha_{V}\right\}$ contained in the region satisfying $\dot{V}=0$, and $\alpha_{V}\leq\gamma(\gamma-\chi_{n}\kappa_{n})/\kappa_{n}^{2}$.
@@
Assume initially $X_{F}\in\Omega_{\alpha}$ and $\Gamma_{F}$ represents an invariant set of $X_{F}$ satisfying that $\dot{V}(X_{F})=0$. When $2\kappa_{n}\alpha_{V}+\gamma\chi_{n}-2\gamma R_{n}<0$, $\dot{V}<0$ is satisfied according to Eqs. (29), (22a) with $\Omega=\tilde{\omega}_{n}$, and $V(x)\geq 0$ by its definition. Then there exists a positive value $v_{l}$ that $\lim_{t\rightarrow\infty}V(X_{F})=v_{l}\geq 0$ and $v_{l}<\alpha_{V}$. Thus $V(X_{F})$ is always within the invariant set. Denote the solutions of $R_{n}$ satisfying $\dot{V}=0$ as $r_{1}$ and $r_{2}$, then $\alpha_{V}\leq r_{1}^{2}+r_{2}^{2}=\gamma(\gamma-\chi_{n}\kappa_{n})/\kappa_{n}^{2}$. ∎
@@
Taking the circumstance $N=2$ as an example, the real and imaginary parts of $F_{1}(t)$ are compared in Fig. 2, where we take $\chi_{1}=1$, $\gamma=2$GHz [55], $\kappa_{1}=1$GHz, $\Omega=50$GHz and $\mathrm{d}t=0.01$ns. The comparisons in Figs. 2(a) and (b) show that, the diverse between $\Omega$ and $\tilde{\omega}_{1}$ can induce oscillations of $R_{1}(t)$ and $I_{1}(t)$, and this can influence the non-Markovian dynamics via the time-varying decay rates in Eq. (9).
@@
II-C Nonlinear dynamics with an uncertain input
@@
According to Eqs. (17)-(20), when $\Omega=\tilde{\omega}_{n}$, $I_{n}(t)\equiv 0$, and we denote $\lim_{t\rightarrow\infty}R_{n}(t)=\bar{R}_{n}$. Define a new state vector $\tilde{X}_{F}(t)=\left[R_{n}(t)-\bar{R}_{n},I_{n}(t)\right]^{\rm T}\triangleq\left[\tilde{R}_{n}(t),I_{n}(t)\right]^{\rm T}$ and $\lim_{t\rightarrow\infty}\tilde{X}_{F}(t)=\left[0,0\right]^{\rm T}$ when $\Omega=\tilde{\omega}_{n}$ and Assumption 1 is satisfied. Then
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{R}_{n}(t)=\kappa_{n}\left[\left(\tilde{R}_{n}(t)+I_{n}\right)^{2}-I_{n}^{2}(t)\right]-\gamma\left(\tilde{R}_{n}(t)+\bar{R}_{n}\right)+\left(\Omega-\tilde{\omega}_{n}\right)I_{n}(t)+\frac{\gamma\chi_{n}}{2},$ (31a)
$\frac{\mathrm{d}}{\mathrm{d}t}I_{n}(t)=2\kappa_{n}\left(\tilde{R}_{n}(t)+\bar{R}_{n}\right)I_{n}(t)-\left(\Omega-\tilde{\omega}_{n}\right)\left(\tilde{R}_{n}(t)+\bar{R}_{n}\right)-\gamma I_{n}(t),$ (31b)
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{X}_{F}(t)=$ $\tilde{\textbf{f}}\left(\tilde{X}_{F}(t)\right)+\begin{bmatrix}\frac{\gamma\chi_{n}}{2}+\kappa_{n}\bar{R}_{n}^{2}-\gamma\bar{R}_{n}\\ \left(\tilde{\omega}_{n}-\Omega\right)\bar{R}_{n}\end{bmatrix}+\begin{bmatrix}0&\Omega-\tilde{\omega}_{n}\\ \tilde{\omega}_{n}-\Omega&0\end{bmatrix}\tilde{X}_{F}(t),$ (32)
@@
When $\tilde{\omega}_{n}=\Omega$, by Theorem 1, the quantum dynamics approaches Markovian behavior as $t\rightarrow\infty,$ and the nonlinear dynamics $\dot{\tilde{X}}_{F}(t)=\tilde{\textbf{f}}\left(\tilde{X}_{F}(t)\right)$ is uniformly and asymptotically stable around its equilibrium $\tilde{X}_{F}=0$. When $\left|\tilde{\omega}_{n}-\Omega\right|\leq\epsilon$ and $\gamma\chi_{n}$ is finite, Proposition 6 in Appendix C ensures that $\tilde{R}_{n}(t)$ and $I_{n}(t)$ in Eq. (31) remain stable. Then the trajectory of $X(t)$ in Eq. (9) converges to be Markovian as $t\rightarrow\infty$. ∎
@@
Now we take $N=3$ as an example for numerical simulations of the above non-Markovian dynamics around the parameter settings in Refs. [56, 57]. As shown in Fig. 3, we take $\omega_{1}=37.7$GHz, $\omega_{2}=75.3$GHz, $\Omega=50$GHz, $\Delta_{1}=-20$MHz, $\Delta_{2}=-70$MHz, $g_{1}=g_{2}=20$MHz, $\chi_{1}=\chi_{2}=1$, $\gamma=10$GHz, $\kappa_{1}=\kappa_{2}=0.31$GHz, and the time step for simulation equals 0.010.01ns. In the simulations for Markovian circumstance, we take $F_{n}$ with $n=1,2$ as constants that equal the steady values of those in the non-Markovian simulations.
Figure 3: Non-Markovian and Markovian dynamics based on the quantum system that one three-level atom is coupled to a cavity.